Metabolizing a candy bar releases about 10^6 joules. How fast must the candy bar travel to have the same 10^6 joules in kinetic energy? The mass of the candy bar is 0.2 kilograms.
1
Expert's answer
2011-05-12T11:21:34-0400
The kinetic energy is
K = mV2/2
Thus
V = √(2*K/m) = √(2*106/0.2) = 2 √(gR) = √(10*39) = 19.75 m/s